OpenStudy (anonymous):

ind a possible formula for the 4th degree polynomial g satsfying the following properties. g has a double zero at X=3; g(5)=0, g(-1)=0, and g(0)=3. You may leave your answer in factored form

OpenStudy (anonymous):

\[g(x)=(x-3)^2(x-b)(x-c)\]

OpenStudy (anonymous):

\[g(x)=a(x-3)^2)(x-5)(x+1)\]

OpenStudy (anonymous):

get a (0,3) \[3=a(0-3)^2(0-5)(0+1)\]

OpenStudy (anonymous):

Is that my ans?

OpenStudy (anonymous):

no \[g(x)=\frac{ 1 }{ 15 }(x-3)^2(x-5)(x+1)\]

OpenStudy (anonymous):

?